# A QUBO Formulation for the Generalized LinkedIn Queens Game

### v1

@@ 7 Conclusions
We have presented several generalizations of the N-queens problem, including the case of LinkedIn, presenting a general QUBO formulation to solve them, and interesting particular cases with their corresponding formulations. We have also presented two new interesting types of problems, associated with N-queens, and their respective QUBO formulations.
Future lines of research may include its implementation in various types of quantum hardware, benchmarking them using these problems, its extension to the case in which we allow more than two consecutive values for the number of queens in each region or its transformation to QUDO formulation.

### v2

@@ 4 Conclusions
We have presented several generalizations of the N-queens problem, including the case of LinkedIn, presenting a general QUBO formulation to solve them, and interesting particular cases with their corresponding formulations, as the Tents & Trees. We have also presented two new interesting types of problems, associated with N-queens, and their respective QUBO formulations. We have also presented a mathematical formulation for the Takuzu and the LinkedIn Tango games, shown a generalization of it, and created the QUBO formulation that solves it. We have seen that this formulation requires at most $NM-N_{I0}-N_{I1}-N_{He}-N_{Hc}-N_{Ve}-N_{Vc}$ variables, being greatly optimized thanks to the initial conditions and equal/cross terms. We have also presented several generalizations of the problem and their corresponding QUBO formulations.
Future lines of research may include its implementation in various types of quantum hardware, benchmarking them using these problems, its extension to the case in which we allow more than two consecutive values for the number of queens in each region, the creation of a Grover’s [14] oracle to solve the problem, its transformation to QUDO or HOBO formulation or addressing how to efficiently enforce the row and column non-repetition constraint.